A can travels for two hours at an average speed of 40 miles per hour. For how long must the can then travel at an average speed of 55 miles per hour to raise the overall average speed to 45 miles per hour? What is the total time traveled?
Answer by scott8148(5895)

let x="time @ 55mph"

(2(40)+x(55))/(2+x)=45

multiplying by 2+x __ 80+55x=90+45x __ 10x=10